# mCLM: A Function-Infused and Synthesis-Friendly Modular Chemical Language Model

### v1

@@ Appendix A Pretraining Data Mixture
To train the model, our goal is to cover a wide variety of different molecular functions and classes of molecules. To do so, we consider a wide variety of data sources. We use three existing instruction datasets: SMolInstruct [115], Tulu-3 [44], and the Biomolecular text portion of Mol-Instructions [24]. SMolInstruct is used to cover the standard set of chemistry LLM tasks. We supplement this with the non-overlapping portion of Mol-Instructions. Finally, we include Tulu-3 to preserve general reasoning capability.
@@
A.1 Property-Based Synthetic Instruction Data Generation for Contrastive Learning
@@
In addition to these tasks, we also consider the molecule optimization task as described in DrugAssist [114] (e.g., given molecule A, improve property X). However, their approach has fundamental limitations due to the reliance on oracle models. These models may be out-of-distribution for molecules within our dataset, and low-quality models may propagate errors into our training data. To address this issue, we consider “activity cliffs” [85, 121] within existing property datasets to train the model for molecule optimization. Activity cliffs are two molecules which are structurally similar but have large differences in a given property. We opt to base our definition of activity cliff on Murcko scaffolds [7]. This methodology, which is widely used in medicinal chemistry, extracts the core subgraph, or scaffold, of a molecule. We found that this approach was much more computationally feasible on large-scale molecule datasets than approaches like fingerprint similarity [6].
@@
Given nominal data (e.g., toxic, kinase inhibitor, banana smell), we look for a pair of molecules that have the same scaffold but only one molecule has the desired property. Given numerical data (e.g., solubility, power conversion efficiency, HOMO-LUMO gap), we instead look for molecules which have a property difference of 1 standard deviation. This forms the positive-negative same-scaffold portion of our data. We also consider pairs of molecules that have different scaffolds but the same property. Here, our goal is to teach the model to consider function over structure. As an example, we may ask “Propose a molecule with a different structure than [MOL] … [/MOL] that still demonstrates anticoagulant properties. [MOL] … [/MOL]”. For a select number of property types, we use oracle models instead of ground truth data. Please see Appendix B.2.3 for more information.
A.2 Data Filtering
@@
Due to the large size of our dataset, we filter our data to only keep the most frequent 100k (50k end blocks and 50k mid blocks), 10k (5k and 5k), and 1000 blocks (500 and 500) for experiments. This is because there are far fewer molecule tokens compared to text (each training sample only has up to about 10). Due to Zipf’s law [127], in the full dataset, a single molecule token may only appear in one example in the entire training corpus. Without considerably scaling the compute budget, understanding this additional block is difficult. Overall, this filtering enables us to learn more efficiently and requires less memory resources for training the model, and allows us to better test the model architecture. Please see the section on training details for more information in Appendix C and see Appendix E for source property datasets.
Data Type Full Data Only Blocks Top 100k Top 10k Top 1000 All 36,641,170 20,910,431 14,823,280 6,195,903 1,054,124 Existing Data Sources 3,117,841 1,208,365 1,049,473 537,674 149,994 SMolInstruct 2,124,738 215,262 56,370 14,242 2,300 Tulu-3 939,343 939,343 939,343 469,672 93,934 Mol-Instructions Biomol. Text 53,760 53,760 53,760 53,760 53,760 Synthetic Real Data 21,290,353 9,294,434 7,711,548 3,960,633 705,210 Classification 5,612,215 3,771,038 3,192,583 1,337,619 223,341 Molecule Generation 535,317 316,630 245,034 150,001 27,393 Positive Negative Same Scaffold 3,795,820 113,390 65,566 16,286 2,007 Positive Positive Different Scaffold 660,925 477,976 256,435 191,720 33,579 Property to Molecule 6,676,957 3,125,097 2,878,012 1,639,322 307,780 Multi-Property to Molecule 572,505 556,107 533,647 423,089 78,793 Scaffold+Property to Molecule 705,833 405,497 226,696 80,448 13,078 Regression 2,730,781 528,699 313,575 122,148 19,239 Synthetic Oracle Data 12,232,976 10,407,632 6,062,259 1,697,596 198,920 Classification 2,077,327 1,778,874 1,058,754 299,444 35,168 Positive Negative Same Scaffold 72,763 58,645 25,035 4,935 469 Scaffold+Property to Molecule 1,169,407 925,055 453,087 113,049 13,642 Property to Molecule 13,257 13,254 8,841 906 2 Regression 8,900,222 7,631,804 4,516,542 1,279,262 149,639
@@
Table 3: Breakdown of molecule data.
Appendix B Synthetic Data Generation
B.1 Task Formulation for Molecular Design
@@
B.2 Instruction Tuning Data Generation
B.2.1 Prompt-Answer Templates
@@
B.2.2 Label Sources for Instruction Tuning Data
@@
B.2.3 Ensemble Oracle Model:
@@
  • FARM [68]: A SMILES-based BERT model trained with functional group-aware tokenization.
  • ChemBERTA-2 [2]: A large-scale transformer model trained on millions of canonical SMILES sequences.
  • GNN [19]: A graph neural network trained on molecular graphs with atom- and bond-level features.
@@
FARM [68] 0.88 0.89 0.79 0.94 0.88 0.92
GNN [19] 0.75 0.78 0.86 0.79 0.80 0.81
ChemBERTa-2 [2] 0.86 0.89 0.81 0.93 0.86 0.99
@@
Appendix C Training Procedure
We employed Qwen2.5-3B [112] as the starting LLM for building the mCLM. Generally, we followed the training procedure from LLaVa [56, 58, 47]. We used a two-layer MLP with PReLU activation [30] as an adapter into the LLM input/output from the GNN. We selected an initial learning rate of 1e-5 for the full model and 1e-6 for the adaptor and LM heads. Further, we used a cosine annealing schedule with a minimum of 1e-6 and 2000 linear warmup steps; AdamW [63] optimizer was employed. The model was trained on 4 A100 80GB GPUs in bfloat16 precision.
We found that the model learned the molecule tokens much slower than the text (there is usually a 10x difference in loss value). Molecule tokens are rarer and show up less in the training data. Because of this, we decided to separate the molecule classifier head and the language classifier head. We used a standard autoregressive language modeling loss for both, and we averaged these two losses for the final loss value. The main part of our training experiments focused on minimizing the molecule loss, since the text loss was easy to optimize. Further, we found PEFT [31] was not sufficient to adapt to molecules, so full finetuning was required. Roughly 10-50 examples from each synthetic (data source, task) pair were put into a validation set.
To initialize the GNN weights, we employed the MolCLR [103] unsupervised contrastive learning technique. We used AugliChem [66] for the augmentations: random atom masking, random bond deletion, and motif removal. One of these augmentation was selected uniformly at random for each data point. The GNN was initialized using a batch size of 128 and lr of 1e-4 with a cosine schedule. The model was trained on all 800k blocks in the full data until convergence on a validation set. Please see Appendix LABEL:appendix:molecule_token_examples for vocabulary examples and their nearest neighbor embeddings. We tested embedding dimensions between 16 and 4,096 and found 128 dimensions to be sufficient while minimizing total memory cost. This was necessary because we stored the entire embedding matrix in GPU memory, which was much faster, but consumed about 20GB VRAM. Doing so allowed us to train without the GNN during our pretraining process, which is considerably more efficient. We note the GNN can then be finetuned along with the rest of the mCLM during finetuning to new types of molecules or specific tasks. While we did consider a sampled softmax to train the mCLM, we found this to limit the learning of the model.
For training the mCLM, we used two stages for pretraining. First, we trained for 1 epoch with everything frozen except the adaptor, to allow the adaptor to adjust to the LLM representation space. For the second stage, we trained for 5 epochs with only the GNN embeddings frozen. As discussed in the training data mixture section A, we used the most frequent 1000 building blocks as our vocabulary.
@@ Appendix C Training Procedure
Appendix D Tokenizer Details
D.1 Synthetic Tokenizer Details
@@
The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a pre-determined small set of reactions, preferably only those that can be performed in an automated manner. For instance, if amide bond formation is defined as available, the tokenizer will be able to disconnect amide bonds in the molecule of interest. Up to this point, the protocol is synonymous with classical computational retrosynthesis but there is a fundamental difference. Namely, the sets of reactions suitable for automated synthesis is very limited – in fact, state-of-the-art synthesis machines utilize only three types of disconnections (amide bond formation, as well as Suzuki and Buchwald-Hartwig couplings). This places very stringent requirements on the groups that can be present in the disconnected blocks – for instance, when the disconnection (say, Buchwald-Hartwig coupling) yields an amine functionality on one of the blocks, this block cannot contain any groups that during the anticipated uses of this block would present a synthetic incompatibility. In the most trivial case, the block cannot contain another unprotected amine because after the Buchwald-Hartwig disconnection, the block would feature two amines which, in turn, would present competing reactive sites (in the Buchwald-Hartwig synthesis but also in the formation of amide bonds). Therefore, the tokenizer performs retrosynthetic operations while simultaneously checking if they do not lead to blocks with functional groups presenting competing reactivities. Only disconnections avoiding such problems are allowed. This then guarantees that when the corresponding blocks are used to make other molecules, they give only the selective synthesis outcomes.
Appendix E Source Datasets and Databases
  • Leffingwell odors [76]
  • BACE [108]
  • Flashpoint [86]
  • MUV [108]
  • Tox21 [108]
  • AMES [33]
  • Bioavailability [33]
  • Caco2 [33]
  • Carcinogens [33]
  • cav3_t [33]
  • choline_transporter [33]
  • clearance hepatocyte [33]
  • CYP1A2 [33]
  • CYP2C9 [33]
  • CYP2D6 [33]
  • CYP3A4 [33]
  • DILI (drug induced liver toxicity) [33]
  • Half_life [33]
  • hERG [33]
  • HIA (human intestinal absorption) [33]
  • Hydration free energy [33]
  • kcnq2 [33]
  • LD50 [33]
  • Lipophilicity [108]
  • m1_muscarinic [33]
  • OPV data [69, 62]
  • orexin_receptor [33]
  • PAMPA_NCATS [33]
  • Broccatelli [9]
  • potassium_ion_channel [33]
  • PPBR (Plasma Protein Binding Rate) [33]
  • Pubchem logP [40]
  • SARSCoV2_3CL [33]
  • SARSCoV2_vitro [33]
  • serine_threonine_kinase_33 [33]
  • Skin Reaction [33]
  • Solubility_AqSolDB [33]
  • tyrosyl-dna_phosphodiesterase [33]
  • VDss (volume of distribution at steady state) [33]
  • Molecule Property Cliff Datasets (30+ datasets) [94]
  • Chemical Function (CheF) [42]
  • ChemFOnt: the chemical functional ontology resource [106]
  • Pubchem properties [40]
  • FreeSolv [108]
  • QM8 [108]
  • QM9 [108]
  • Thermosol [108]
  • ESOL (Estimated SOLubility) [108]
  • Lipo [108]
  • BBBP [27]
  • ClinTox [108]
  • HIV [108]
  • SIDER [108]
  • Forward synthesis (USPTO) [115]
  • Retrosynthesis [115]
  • CheBI-20 [16]
  • L+M-24 [21]
  • HBA [27]
  • HBD [27]
  • MolWt [27]
  • NumAromaticRings [27]
  • rotatable_bonds [27]
  • TPSA (topological polar surface area) [27]

### v2

@@ Appendix C Pretraining Data Mixture
To train the model, our goal is to cover a wide variety of different molecular functions and classes of molecules. To do so, we consider a wide variety of data sources. We use three existing instruction datasets: SMolInstruct [Yu et al., 2024], Tulu-3 [Lambert et al., 2024], and the Biomolecular text portion of Mol-Instructions [Fang et al., 2023]. SMolInstruct is used to cover the standard set of chemistry LLM tasks. We supplement this with the non-overlapping portion of Mol-Instructions. Finally, we include Tulu-3 to preserve general reasoning capability.
@@
C.1 Property-Based Synthetic Instruction Data Generation for Contrastive Learning
@@
In addition to these tasks, we also consider the molecule optimization task as described in DrugAssist [Ye et al., 2025] (e.g., given molecule A, improve property X). However, their approach has fundamental limitations due to the reliance on oracle models. These models may be out-of-distribution for molecules within our dataset, and low-quality models may propagate errors into our training data. To address this issue, we consider “activity cliffs” [Stumpfe et al., 2019, Zhang et al., 2023] within existing property datasets to train the model for molecule optimization. Activity cliffs are two molecules which are structurally similar but have large differences in a given property. We opt to base our definition of activity cliff on Murcko scaffolds [Bemis and Murcko, 1996]. This methodology, which is widely used in medicinal chemistry, extracts the core subgraph, or scaffold, of a molecule. We found that this approach was much more computationally feasible on large-scale molecule datasets than approaches like fingerprint similarity [Bajusz et al., 2015].
Figure 7: Overview of Data Creation.
@@
Given nominal data (e.g., toxic, kinase inhibitor, banana smell), we look for a pair of molecules that have the same scaffold but only one molecule has the desired property. Given numerical data (e.g., solubility, power conversion efficiency, HOMO-LUMO gap), we instead look for molecules which have a property difference of 1 standard deviation. This forms the positive-negative same-scaffold portion of our data. We also consider pairs of molecules that have different scaffolds but the same property. Here, our goal is to teach the model to consider function over structure. As an example, we may ask “Propose a molecule with a different structure than [MOL] … [/MOL] that still demonstrates anticoagulant properties. [MOL] … [/MOL]”. For a select number of property types, we use oracle models instead of ground truth data. Please see Appendix D.2.3 for more information.
C.2 Data Filtering
@@
Due to the large size of our dataset, we filter our data to only keep the most frequent 100k (50k end blocks and 50k mid blocks), 10k (5k and 5k), and 1000 blocks (500 and 500) for experiments. This is because there are far fewer molecule tokens compared to text (each training sample only has up to about 10). Due to Zipf’s law [Zipf, 1936], in the full dataset, a single molecule token may only appear in one example in the entire training corpus. Without considerably scaling the compute budget, understanding this additional block is difficult. Overall, this filtering enables us to learn more efficiently and requires less memory resources for training the model, and allows us to better test the model architecture. Please see the section on training details for more information in Appendix E and see Appendix G for source property datasets. Table 5 gives a breakdown data quantity by instruction type for each subset of the data. Table 6 gives a breakdown of the molecule data for each subset of the data. Note that not all molecules in the “Full Data” can be represented with blocks, so we have a subset for “Only Blocks” molecules, and then molecules consisting of the most frequent “Top 100k” blocks, “Top 10k” blocks, and “Top 1000” blocks. Scaffolds are the number of unique Murcko scaffolds [Bemis and Murcko, 1996]. “Caps” are end blocks, while “Mid” are middle blocks. “Synthesis-aware” is the number of synthesis-guaranteed blocks out of “Total”. Data splits are subdivided into their original source: our constructed instruction data or molecule instructions from SMolInstruct.
@@
Data Type Full Data Only Blocks Top 100k Top 10k Top 1000 All 36,641,170 20,910,431 14,823,280 6,195,903 1,054,124 Existing Data Sources 3,117,841 1,208,365 1,049,473 537,674 149,994 SMolInstruct 2,124,738 215,262 56,370 14,242 2,300 Tulu-3 939,343 939,343 939,343 469,672 93,934 Mol-Instructions Biomol. Text 53,760 53,760 53,760 53,760 53,760 Synthetic Real Data (Ours) 21,290,353 9,294,434 7,711,548 3,960,633 705,210 Classification 5,612,215 3,771,038 3,192,583 1,337,619 223,341 Molecule Generation 535,317 316,630 245,034 150,001 27,393 Positive Negative Same Scaffold 3,795,820 113,390 65,566 16,286 2,007 Positive Positive Different Scaffold 660,925 477,976 256,435 191,720 33,579 Property to Molecule 6,676,957 3,125,097 2,878,012 1,639,322 307,780 Multi-Property to Molecule 572,505 556,107 533,647 423,089 78,793 Scaffold+Property to Molecule 705,833 405,497 226,696 80,448 13,078 Regression 2,730,781 528,699 313,575 122,148 19,239 Synthetic Oracle Data (Ours) 12,232,976 10,407,632 6,062,259 1,697,596 198,920 Classification 2,077,327 1,778,874 1,058,754 299,444 35,168 Positive Negative Same Scaffold 72,763 58,645 25,035 4,935 469 Scaffold+Property to Molecule 1,169,407 925,055 453,087 113,049 13,642 Property to Molecule 13,257 13,254 8,841 906 2 Regression 8,900,222 7,631,804 4,516,542 1,279,262 149,639
Table 6: Breakdown of molecule data by dataset category.
@@
Appendix D Synthetic Data Generation
D.1 Task Formulation for Molecular Design
@@
D.2 Instruction Tuning Data Generation
D.2.1 Prompt-Answer Templates
@@
D.2.2 Label Sources for Instruction Tuning Data
@@
D.2.3 Ensemble Oracle Model:
@@
  • FARM [Nguyen et al., 2024a]: A SMILES-based BERT model trained with functional group-aware tokenization.
  • ChemBERTA-2 [Ahmad et al., 2022]: A large-scale transformer model trained on millions of canonical SMILES sequences.
  • GNN [Edwards et al., 2024a]: A graph neural network trained on molecular graphs with atom- and bond-level features.
@@
FARM [Nguyen et al., 2024a] 0.88 0.89 0.79 0.94 0.88 0.92
GNN [Edwards et al., 2024a] 0.75 0.78 0.86 0.79 0.80 0.81
ChemBERTa-2 [Ahmad et al., 2022] 0.86 0.89 0.81 0.93 0.86 0.99
@@
Appendix E Training Procedure
We employed Qwen2.5-3B [Yang et al., 2024] as the starting LLM for building the mCLM. Generally, we followed the training procedure from LLaVa [Liu et al., 2023a, 2024b, Li et al., 2024a]. We used a two-layer MLP with PReLU activation [He et al., 2015] as an adapter into the LLM input/output from the GNN. We selected an initial learning rate of 1e-5 for the full model and 1e-6 for the adaptor and LM heads. Further, we used a cosine annealing schedule with a minimum of 1e-6 and 2000 linear warmup steps; AdamW [Loshchilov and Hutter, 2017] optimizer was employed. The model was trained on 4 A100 80GB GPUs in bfloat16 precision.
We found that the model learned the molecule tokens much slower than the text (there is usually a 10x difference in loss value). Molecule tokens are rarer and show up less in the training data. Because of this, we decided to separate the molecule classifier head and the language classifier head. We used a standard autoregressive language modeling loss for both, and we averaged these two losses for the final loss value. The main part of our training experiments focused on minimizing the molecule loss, since the text loss was easy to optimize. Further, we found PEFT [Hu et al., 2021] was not sufficient to adapt to molecules, so full finetuning was required. Roughly 10-50 examples from each synthetic (data source, task) pair were put into a validation set.
To initialize the GNN weights, we employed the MolCLR [Wang et al., 2022b] unsupervised contrastive learning technique. We used AugliChem [Magar et al., 2022] for the augmentations: random atom masking, random bond deletion, and motif removal. One of these augmentation was selected uniformly at random for each data point. The GNN was initialized using a batch size of 128 and lr of 1e-4 with a cosine schedule. The model was trained on all 800k blocks in the full data until convergence on a validation set. We tested embedding dimensions between 16 and 4,096 and found 128 dimensions to be sufficient while minimizing total memory cost. This was necessary because we stored the entire embedding matrix in GPU memory, which was much faster, but consumed about 20GB VRAM. Doing so allowed us to train without the GNN during our pretraining process, which is considerably more efficient. We note the GNN can then be finetuned along with the rest of the mCLM during finetuning to new types of molecules or specific tasks. While we did consider a sampled softmax to train the mCLM, we found this to limit the learning of the model.
For training the mCLM, we used two stages for pretraining. First, we trained for 1 epoch with everything frozen except the adaptor, to allow the adaptor to adjust to the LLM representation space. For the second stage, we trained for 5 epochs with only the GNN embeddings frozen. As discussed in the training data mixture section C, we used the most frequent 1000 building blocks as our vocabulary.
@@ Appendix E Training Procedure
Appendix F Tokenizer Details
F.1 Synthesis-Guaranteed Tokenizer Details
@@
The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a pre-determined small set of reactions, preferably only those that can be performed in an automated manner. For instance, if amide bond formation is defined as available, the tokenizer will be able to disconnect amide bonds in the molecule of interest. Up to this point, the protocol is synonymous with classical computational retrosynthesis, but there is a fundamental difference. Namely, the sets of reactions suitable for automated synthesis is very limited – in fact, state-of-the-art synthesis machines utilize only three types of disconnections (amide bond formation, as well as Suzuki and Buchwald-Hartwig couplings). This places very stringent requirements on the groups that can be present in the disconnected blocks – for instance, when the disconnection (say, Buchwald-Hartwig coupling) yields an amine functionality on one of the blocks, this block cannot contain any groups that during the anticipated uses of this block would present a synthetic incompatibility. In the most trivial case, the block cannot contain another unprotected amine because after the Buchwald-Hartwig disconnection, the block would feature two amines which, in turn, would present competing reactive sites (in the Buchwald-Hartwig synthesis but also in the formation of amide bonds). Therefore, the tokenizer performs retrosynthetic operations while simultaneously checking if they do not lead to blocks with functional groups presenting competing reactivities. Only disconnections avoiding such problems are allowed. This then guarantees that when the corresponding blocks are used to make other molecules, they give only the selective synthesis outcomes.
F.2 Molecule Token Examples
As discussed in E, we pretrained our GNN encoder using MolCLR [Wang et al., 2022b] on all 8̃00k blocks in our full pretraining dataset. Figures 8-13 shows six examples of random blocks and their 5 nearest neighbors in our dataset. The similarity of the blocks and their neighbors shows the high quality representations which were obtained.
Figure 8: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Figure 9: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Figure 10: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Figure 11: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Figure 12: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Figure 13: Example token from pretraining dataset. Similarity above each molecule is ECFP4 Fingerprint Tanimoto similarity from the ground truth molecule.
Appendix G Source Datasets and Databases
  • Leffingwell odors [Sanchez-Lengeling et al., 2019]
  • BACE [Wu et al., 2018]
  • Flashpoint [Sun et al., 2020]
  • MUV [Wu et al., 2018]
  • Tox21 [Wu et al., 2018]
  • AMES [Huang et al., 2021]
  • Bioavailability [Huang et al., 2021]
  • Caco2 [Huang et al., 2021]
  • Carcinogens [Huang et al., 2021]
  • cav3_t [Huang et al., 2021]
  • choline_transporter [Huang et al., 2021]
  • clearance hepatocyte [Huang et al., 2021]
  • CYP1A2 [Huang et al., 2021]
  • CYP2C9 [Huang et al., 2021]
  • CYP2D6 [Huang et al., 2021]
  • CYP3A4 [Huang et al., 2021]
  • DILI (drug induced liver toxicity) [Huang et al., 2021]
  • Half_life [Huang et al., 2021]
  • hERG [Huang et al., 2021]
  • HIA (human intestinal absorption) [Huang et al., 2021]
  • Hydration free energy [Huang et al., 2021]
  • kcnq2 [Huang et al., 2021]
  • LD50 [Huang et al., 2021]
  • Lipophilicity [Wu et al., 2018]
  • m1_muscarinic [Huang et al., 2021]
  • OPV data [Nguyen et al., 2024b, Lopez et al., 2016]
  • orexin_receptor [Huang et al., 2021]
  • PAMPA_NCATS [Huang et al., 2021]
  • Broccatelli [Broccatelli et al., 2022]
  • potassium_ion_channel [Huang et al., 2021]
  • PPBR (Plasma Protein Binding Rate) [Huang et al., 2021]
  • Pubchem logP [Kim et al., 2023]
  • SARSCoV2_3CL [Huang et al., 2021]
  • SARSCoV2_vitro [Huang et al., 2021]
  • serine_threonine_kinase_33 [Huang et al., 2021]
  • Skin Reaction [Huang et al., 2021]
  • Solubility_AqSolDB [Huang et al., 2021]
  • tyrosyl-dna_phosphodiesterase [Huang et al., 2021]
  • VDss (volume of distribution at steady state) [Huang et al., 2021]
  • Molecule Property Cliff Datasets (30+ datasets) [Van Tilborg et al., 2022]
  • Chemical Function (CheF) [Kosonocky et al., 2023]
  • ChemFOnt: the chemical functional ontology resource [Wishart et al., 2023]
  • Pubchem properties [Kim et al., 2023]
  • FreeSolv [Wu et al., 2018]
  • QM8 [Wu et al., 2018]
  • QM9 [Wu et al., 2018]
  • Thermosol [Wu et al., 2018]
  • ESOL (Estimated SOLubility) [Wu et al., 2018]
  • Lipo [Wu et al., 2018]
  • BBBP [Gaulton et al., 2012]
  • ClinTox [Wu et al., 2018]
  • HIV [Wu et al., 2018]
  • SIDER [Wu et al., 2018]
  • Forward synthesis (USPTO) [Yu et al., 2024]
  • Retrosynthesis [Yu et al., 2024]
  • CheBI-20 [Edwards et al., 2021]
  • L+M-24 [Edwards et al., 2024c]
  • HBA [Gaulton et al., 2012]
  • HBD [Gaulton et al., 2012]
  • MolWt [Gaulton et al., 2012]
  • NumAromaticRings [Gaulton et al., 2012]
  • rotatable_bonds [Gaulton et al., 2012]
  • TPSA (topological polar surface area) [Gaulton et al., 2012]